SM Predictions for Top Quarks

The asymmetry in e^{+} e^{−} → bb̄ scattering implies that b is part of SU(2) doublet

• top quark has standard SU(2) couplings

• CKM matrix predicts dominate decay t → W + b

W^{±} and Z^{∘} masses depend on heaviests fermions

• LEP, SLD measurements on the Z resonance require

Mtop = 179 ± 9{+ 17 \atop −19} GeV/c2

D. Schaile, 1995 pp̄ Topical Conference

Top Quark Phenomenology

Top Quarks decay via the weak force

• Predict ∼ 100% t → W + b

• Lifetime depends on mass

Γt = \frac{GF M{3 \atop top}}{8√2π} (1 − \frac{M2 W}{M{2 \atop top}} )2 (1 + 2\frac{M2 W}{M{2 \atop top}} )

• For heavy top (Mtop > ∼ 100 − 120 GeV/c2), decays before it hadronises

Sensitive to some classes of “new physics”

• Higgs couples strongly to top

• Could decay into a SUSY H+

t → H+ b

|_{⟶}τ+ ντ

What Do We Already Know About Top?

Searches at e^{+} e^{−} Colliders have ruled out “light” top

• data from LEP requires Mtop > 46 GeV/c2 at 95% CL • Independent of how top decays

Width of W ± boson rules out W + → tb̄

• Mtop > 62 GeV/c2 at 95% CL

F.Abe et al. (CDF), PRL 73, 220 (1994).

Searches at pp̄ colliders at CERN and Fermilab

• searched for a “Standard Model” top

• Mtop > 131 GeV/c2 at 95% CL (D0 Collab).

S. Abachi et al. (D0 / ), PRL 72, 2138 (1994).

Production of Top in Hadron Collisions

Top quark production is mediated by

qq̄, gg → tt̄
qq̄ → tb̄X  and  qq̄ → W + → tb̄ both small

Because top is heavy, need large x for initial partons

• Implies that cross section falls with mass

• t t system approximately at rest – < PT >∼ Mtop/2

Higher order corrections to σ_{t t} have been calculated

• Most recently by Laenen, Smith and Neerven, NP B369, (1992) 543.

• Biggest effects due to “soft gluon emission”

Top Decays and Detection Requirements

SM top quark decays via the weak force

t → W + b

|_{⟶}ud̄ or cs̄ (2/3)

|_{⟶}e+ νe or μ+ νμ or τ+ ντ (1/3)

In proton-antiproton collisions,

This results in several distinct decay signatures

• Six jet final states (36/81)

• Lepton (e± or μ±) + Jet events (24/81)

• Dilepton events with 2 energetic e± and μ∓ (4/81)

• Decay modes with τ’s hard! (∼ 12/81)

Detector Requirements

Detection of these decay modes requires

• Good jet reconstruction (energy and direction)

• Good lepton ID capability (acceptance and efficiency) – > 80 − 85% is reasonable

• Missing energy detection (for neutrinos)

Production/decay properties imply that

• Production is “central” (especially for high mass)

• Most likely initial state is where x1 ∼ x2

• Top quarks are produced with modest PT

⇒ daughters are isotropic in lab frame

Detector requirements boil down to

1. Coverage over |η| < ∼ 2 − 2.5 desirable

2. Good efficiency for partons with PT ≥ 15 − 20 GeV/c 3. Good lepton ID with high efficiency

Can see this from t t MC

• Mtop = 160 GeV/c2 in dilepton channel • Plot observed jet and lepton energies

Background Considerations

The backgrounds are dominated by QCD processes

• Multi-jet final states direct background to 6-jet channel • Jets faking leptons potential background to lepton ID • Many SM EWK physics processes

– Drell-Yan production in e+ e− and μ+ μ− modes

– Z∘ production

– Semi-leptonic bottom and charm quark decays

– QCD W + jet processes

– Di-vector boson final states: W + W −, W + Z∘

Any of these could compromise one of the searches

• Lepton ID must have extremely good jet-rejection

• Kinematics of background events must be used

– For example, can use PT and lepton isolation requirements to reduce and characterise b, c quark production

Background Studies Extremely Important

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

CDF performs search requiring

• Two clean, high PT lepton candidates

• Missing transverse energy E̸T > 25 GeV

• Require two jets with ET > 10 GeV and |η| < 2.0 • Reject specific backgrounds (e.g. Z∘)

Significant dilepton background is Drell-Yan production

• E̸T signature created by mismeasured jet

• Harden E̸T cut to > 50 GeV if correlated with jet

a) Z+1 jet, b) Z+ 2 jet, and c) t t (Mtop = 160 GeV/c2)

Example: CDF Z → τ^{+} τ^{−} Background

The Z → τ^{+} τ^{−} → e^{+} μ^{−}ν_{e}ν̄_{μ}X is largest background for all dilepton channels

• Take sample of observed Z → e+ e− and Z → μ+ μ− events – Gives correct Z kinematics

– Gives correct associated jet production

• Replace each lepton with a τ± – Use MC to decay τ±

– Use detector simulation on τ± decay daughters

Resulting calculation is relatively model-independent

• MC model dependence very small

• Limited by Z statistics

Jet Cut  e^{+} e^{−} and μ^{+} μ^{−}  e^{±}μ^{∓}
Jet E_{T} > 10 GeV  0.17 ± 0.05  0.21 ± 0.05
Jet E_{T} > 15 GeV  0.10 ± 0.03  0.13 ± 0.04

Results for Dilepton Channel

Have measured the rate of events from non-top sources

• Used data to measure most of these rates

• Cross-checked using Z∘ → e+ e− and Z∘ → μ+ μ−

• Consistent with theoretical calculations

Background  eμ (events)  ee, μμ (events)
Z → τ+ τ−  0.21 ± 0.05  0.17 ± 0.05
One or both leptons fake  0.10 ± 0.10  0.13 ± 0.10
bb̄  0.01 ± 0.01  0.02 ± 0.01
W W, W Z and ZZ  0.13 ± 0.06  0.08 ± 0.03
Drell-Yan  0.44 ± 0.28
Total  0.45 ± 0.14  0.84 ± 0.28

• Expect 1.3 ± 0.3 background events

Control of Backgrounds

Evolution of background rates in dileptons instructive

Run  Sensitivity  Bkg σ Comments
Period  (pb−1)  (fb)
1988-89  4  350  Just dilepton requirement
CDF 1A  20  20  More Drell-Yan rejection
CDF Run 1A/B  67  19  Added ≥ 2 jet cut
D0 / 1A  15  72  ≥ 1 jet (for Mtop< 120)
D0 / 1A  15  52  ≥ 2 jets (for Mtop> 120)
D0 / 1A/B  44-56  14  Added HT cut

Have to be careful with background studies

• Data fluctuations can lead to “carving up” of acceptance

• Base cuts on independent samples

• Use MC predictions to guide

What Is Expected From Top

Expected number of t t→ dileptons is determined using MC and full detector simulation

Mtop  σt t  Events
(GeV/c2) (pb) Detected
160  8.2  4.4
170  5.8  3.0
180  4.2  2.4

Breakdown by channel reflects branching ratios

Channel Signal Fraction
eμ  57%
ee  20%
μμ  23%

Mtop of 170 GeV/c2 assumed

Dilepton acceptance as a function of top mass

15

Dielectron Data for Run 1A and 1B

16

DiMuon Data for Run 1A and 1B

17

Electron-Muon Data for Run 1A and 1B

18

Summary of CDF Dilepton Search

Observe a total of 6 candidate events

• Observe 1 μ+ μ− event and 5 eμ events – Additional μ+ μ− event excluded

– Has μ+ μ−γ mass close to Z∘

• Expect 1.3 ± 0.3 events from background

Probability of background fluctuation is 3.3 × 10^{−3}

This means:

• “If experiment is repeated and top doesn’t exist, then fraction of experiments with at least 6 events observed is 1 in 300.”

B Quark Tags in CDF Dilepton Events

Can a posteriori check for evidence of b jets

• Use SVX and SLT algorithms described later – Two events with SVX-SLT “double tags”

– One event with SLT μ tag

• Would expect 0.47 tags if no b quark jets – Would expect 3.6 tags if all 6 are t t

Lepton + Jets Search Philosophy

Decay mode with the greatest sensitivity is “lepton+jets”

• About 30% of t t events decay this way

• Look for one high PT lepton

• Require E̸T > 25 GeV to “tag” neutrino

Problems come from physical backgrounds

• Backgrounds from “fake” leptons, etc. are manageable

• “QCD W + jet” production much larger than t t

• Have to find additional rejection power 1. Place more stringent kinematical requirements 2. Require presence of b quark jet

– “tag” semileptonic b quark decays

– “tag” using large cτ of b decays

∗ This capability is unique to CDF

• Have to maintain good top efficiency

W+Jet Selection

Selection of W candidates straightforward

• Require high PT isolated lepton candidate

• Require significant E̸T

• Require three or more high PT jets

One obtains a large inclusive W samples

• CDF has 31,000 eνe and 21,000 μνμ candidates

Run 1A CDF data shows clear “W” signal:

W + Jet Sample

Require ≥ 3 jets (from other W and b quarks)

• ≥ 3 jets with ET > 15 GeV and |η| < 2.0

– finds 203 W+≥ 3 jet events

Most background sources are small:

Background  Number of Events (CDF)
Fake leptons and bb̄  18 ± 9
W W, W Z and ZZ  9 ± 4
Z → ττ, ee, μμ  15 ± 3
Total  42 ± 10

QCD W+jets production is largest expected background

Mtop (GeV/c2) Number of Events (CDF)
120  136 ± 21
140  77 ± 10
160  42 ± 7
180  23 ± 2.4
W+jets  175{+56 \atop −28}

Jet Multiplicity

Following plot illustrates the serious S/N problem

• Event distribution versus minimum jet multiplicity • Require jet ET > 15 GeV (corrected)

B Tagging Techniques

Use two properties of b quarks to identify them

1. About 20% of b’s decay via

b → e−νeX and b → μ−νμX

• Look for jets with “soft” leptons in them (SLT tags)

• Expect about 15% of all t t events to have SLT

2. “Long” b lifetime – cτ ∼ 450 microns

• Use Silicon Vertex Detector to identify jets with a secondary vertex

• We call these “SVX” tags

• Expect 42% of t t events to have at least one SVX tag

Idea is to just “count” number of tagged events

SVX B Tags

SVX tag requirements are

• at least 3 tracks with impact parameter d > 3σ – use 2 tracks with tighter cuts in 2nd pass

• track PT > 2 GeV/c

• the tracks are consistent with coming from a common vertex

Used control samples to measure efficiency, fake rate

• inclusive electrons enriched in b → e+ νeX decays

• B → J/ψX, J/ψK±, J/ψK∗ decays

• “generic” jet sample

SVX B Tagging

Can demonstrate that tags are really finding b’s

• Inclusive e±’s are ∼ 40 − 60% from b → e−νec • Look at jet “away” from e± candidate

• See expected number of tags

SVX B Tagging Backgrounds

Can use rate of tags in QCD jet data to estimate background

• Multiply number of jets by per-jet tagging rate ft – ft ∼ 0.01 and is independent of jet pT

• This includes real pp̄ → bX

• Include MC estimate for W c production

• Does provide a conservative background estimate

More rigorous to separate background sources

• Fake tags estimated from rate of −ve decay length tags – Turns out to be ∼ 0.003 tags per jet

• Tags from real heavy flavour (W bb̄ and W cc̄)

– Use QCD matrix element for heavy flavour content ∗ Use HERWIG and calculation by M. Mangano ∗ CDF assigns a 100% uncertainty to rate

– Use known tagging efficiencies

– Calibrate against W + 1 and W + 2 jet data

Method I predicts larger background than Method II

• Use Method II as more reliable background estimate

Results for SVX Background Calculation

Can compare CDF results for Run 1B

• Summarise major sources of background below

• Estimates for CDF Run 1B data only

• Note that W c background is added to both

W + 1 jet  W + 2 jet  W + 3 jet  W + ≥ 4 jet
Method 1  59 ± 7  20 ± 2  5.0 ± 0.7  1.6 ± 0.2
Fakes  10 ± 2  3.5 ± 0.7  0.9 ± 0.2  0.4 ± 0.1
W bb̄ W cc̄  11 ± 9  5.4 ± 4.3  1.4 ± 1.1  0.4 ± 0.3
W c  12 ± 4  3.3 ± 1.0  0.7 ± 0.3  0.2 ± 0.1
Method 2  38 ± 10  15 ± 5  3.7 ± 1.2  1.1 ± 0.3

b Quark Tagging in W+Jet Events

Applied b tagging techniques to W+jet events

• The W + 1 and W + 2 jet samples are ultimate control – Don’t expect significant top contribution

– Theory predicts low b hadron content

Sample  SVX bkg SVX tags  SLT bkg  SLT tags
W+1 jet  50 ± 12  40  159 ± 25  163
W+2 jet  21 ± 7  34  46 ± 7  55
W+≥ 3 jet  6.7 ± 2.1  27  15.4 ± 2.3  23

Summary of Counting Experiment

CDF sees significant excesses in all three channels

Channel Background  Tags Events  Prob
SLT  15.4 ± 2.0  23  22  6 × 10−2
SVX  6.7 ± 2.1  27  21  2 × 10−5
DIL  1.3 ± 0.3  6  3 × 10−3

Probability of background = 1 × 10^{−6}

Probability calculation performed by assuming no t t

• “simulate” large number of experiments

• Count how many experiments would “see” more tags

• Properly takes into account – Uncertainties in background estimates

– Correlations in uncertainties

– Poisson fluctuations in total rate of events